OpenStudy (anonymous):

is it always possible to find expression for the inverse of a function ? (im talking only about the expression and not failure cause of domain)

OpenStudy (anonymous):

no. especially since it is not always possible to find the expression for a function

OpenStudy (anonymous):

well when i have something like :x - (2/x) = (1/y) + 1 can i make x(y) ?

OpenStudy (anonymous):

and for 2y = x^2 - (1/x^2) ?

OpenStudy (anonymous):

cause i could not find..

OpenStudy (amistre64):

the OS seems a bit buggy lately

OpenStudy (anonymous):

are you asking "can i solve for x?"

OpenStudy (anonymous):

yes .. i want x as a function of y

OpenStudy (anonymous):

no i don't think you can in this case because if you think of y as a function of x it is not one to one. so you will not be able to use algebra to solve for x

OpenStudy (anonymous):

you know i want to break something.. i wasted much time for this two .. cause i believed (since it was given in my home work and said : "find the inverse") that i actually can find an expression for it

OpenStudy (anonymous):

the equations that i wrote you is my end point ..

OpenStudy (anonymous):

ok what was the original function?

OpenStudy (anonymous):

for example the real one they gave me (before i got where i got ) is :x/(x^2-x-2) but this is only one of the two that i wrote

OpenStudy (anonymous):

so you have \[y=\frac{x}{x^2-x-2}\] right? and you want the inverse. first off lets graph http://www.wolframalpha.com/input/?i=y%3Dx%2F%28x^2-x-2%29 and see that it is not one to one

OpenStudy (anonymous):

then we could still try to solve, we just wont get a function.

OpenStudy (anonymous):

i dont care.. i just need an expression but i cant find any..

OpenStudy (anonymous):

ok so we start with \[y=\frac{x}{x^2-x-2}\] and solve for x. first get \[yx^2-yx-2y=x\] then \[yx^2-yx-x-2y=0\] and then \[yx^2-(y+1)x-2y=0\] then use the quadratic formula to get x wtih \[a=y, b=-(x+1), c = -2y\]

OpenStudy (anonymous):

grr i cant believe !!

OpenStudy (anonymous):

lol thank you!